OpenStudy (anonymous):

graph the equation by plotting 3 points 2y=3x+11

OpenStudy (anonymous):

To do this you need to find 3 points: To find three points choose values of x and then find the corresponding y value: i.e. let x = 0, 2y = 3(0) +11 y = 11/2 = 5.5 point is (0,5.5) let x = 1 2y = 3(1)+11 2y = 14 y = 7 point is (1, 7) let x = 3 2y = 3(3)+11 2y = 20 y = 10 point is (3,10)

OpenStudy (anonymous):

it cant be a 5.5 it has to be a whole number

OpenStudy (anonymous):

In that case choose a different x,let x - 5 2y = 3(5) +11 2y = 26 y = 13 (5,15)

OpenStudy (anonymous):

can u help me with another